OpenStudy (anonymous):

Which of the following masses is the smallest? 0.006 kilograms 500 milligrams 40 grams 4.2 decigrams

OpenStudy (anonymous):

which 1 do u think is the smallest?

OpenStudy (anonymous):

the miligram

OpenStudy (anonymous):

milli is 10^-3 nd deci is 10^-1

OpenStudy (lgbasallote):

turn them all into grams...that will help 0.006 kilograms = 6 grams 500 miligrams = 0.5 grams 40 grams = 40 grams 4.2 decigrams = 0.42 grams now does that help?

OpenStudy (anonymous):

yes that helps basallote but should i multiply or divide?

OpenStudy (lgbasallote):

what do you mean? to convert them?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

are you familiar with using ratio and proportion? i can introduce a way to convert using that..

OpenStudy (anonymous):

yes i am. go ahead a show me please

OpenStudy (lgbasallote):

a) 1 kilogram = 1000 grams to convert 0.006 kilogram into grams... 1 : 1000 = 0.006 : x multiply the means and extremes.. (1)(x) = (0.006)(1000) x = 6 grams b) 1000 miligram = 1 gram to convert 500 miligrams into grams... 1000 : 1 = 500 : x 500 = 1000x 500/1000 = x 1/2 = x x = 0.5 grams

OpenStudy (lgbasallote):

do i still need to show c and d?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

i dont really get where u got 0.006 O_O

OpenStudy (lgbasallote):

it's the given... 0.006 kilograms

OpenStudy (anonymous):

oh nvm blond moment x.x

OpenStudy (lgbasallote):

hah :))

OpenStudy (lgbasallote):

for c..it's already in grams so no need to convert d)10 decigram = 1 gram 10 : 1 = 4.2 : x 4.2 = 10x 4.2/10 = x x = 0.42 grams

OpenStudy (anonymous):

lol when do i know when to multiply then divide?

OpenStudy (lgbasallote):

look at the ratio and proportion...just multiply the insides then multiply the outsides

OpenStudy (anonymous):

hmm...i dont understand that...

OpenStudy (lgbasallote):

or....if it's bigger..multiply if it's smaller..divide for example 0.006 kilogram kilogram is BIGGER than grams therefore multiply kilogram by 1000 0.006 x 1000 = 6

OpenStudy (lgbasallote):

500 miligrams.. miligrams is smaller than grams therefore divide by 1000 500/1000 = 0.5

OpenStudy (anonymous):

ohh ok. so what if it was asking me grams to miligrams i would...divide?

OpenStudy (anonymous):

no multiply?

OpenStudy (lgbasallote):

oops yeah multiply sorry :p

OpenStudy (lgbasallote):

i got confused too lol

OpenStudy (anonymous):

Lol ohk. Thanks I understand :D

OpenStudy (lgbasallote):

great!!! ^_^